# Subsystem fidelity in two-dimensional conformal field theories

### v1

@@ 3.3 Case III
No higher-order analytical results exist in the literature for these cases, so we compare our predictions with numerical results in the spin-1/2 XX chain. In the XX chain, the above CFT results are expected to apply in the limit $1\ll\ell\ll L$. To match CFT and spin chain results, we require high-efficiency and high-precision numerical evaluation of subsystem fidelity in the spin chain. We calculate subsystem fidelity in the XX chain using the truncation method with truncation number $t=8$ from [43] and show results in Figure 1. We define the subtracted fidelity
@@
which is removed from both CFT and spin chain results for clearer comparison. The figure uses $(\alpha,\bar{\alpha})=(1,0)$, showing good agreement between CFT predictions and spin chain results. We also calculated cases with other values of $(\alpha,\bar{\alpha})$, though the agreement is not always as good, possibly due to finite-size effects or limitations of the numerical method.
Figure 1: Comparison between CFT predictions (solid red lines) and $X$ spin chain results (symbols) for the subtracted fidelity $F_{(2)}-F$. The good agreement validates the short-interval expansion approach.
@@ 4.1 Case I
Using the same factorization formula (3.1) as in the bosonic case, we reproduce the short interval expansion of these results.
@@ 4.2 Case II
Using the expansion organized around the energy operator $\varepsilon$
$F(\rho_{A},\sigma_{A})=1+\ell^{2}\mathcal{F}^{(1/2)}{\varepsilon\varepsilon}+\ell^{4}\big(\mathcal{F}^{(1/2)}{\varepsilon\varepsilon\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{T\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{\bar{T}\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{TT}+\mathcal{F}^{(1/2)}{\bar{T}\bar{T}}\big)+O(\ell^{6}),$ (4.4)
@@ 4.3 Case III
We compare these predictions with numerical results in the critical Ising chain, obtained using the truncation method from [43] with truncation number $t=8$, as shown in Figure 2. We define the subtracted fidelity
@@
This work provides direct evidence for the perturbative distinguishability between black hole microstates and the canonical ensemble thermal state, and this contrasts with the indirect evidence from earlier studies [47, 28, 30, 31, 32]
@@ Acknowledgements
J.Z. is grateful to the organizers and participants of the Annual Academic Conference of the Gravitation and Relativistic Astrophysics Division of the Chinese Physical Society (Kunming, April 2025), the Symposium on Gravity and Cosmology (Xi’an, June 2025), and the Symposium on Strings, Gravity, and Gravitational Waves (Rizhao, August 2025), where preliminary results of this work were presented and benefitted from valuable discussions. J.Z. also wishes to thank the Institute of Fundamental Physics and Quantum Technology at Ningbo University for its hospitality during a productive visit where part of this research was presented and completed. In addition, J.Z. thanks Bin Chen, Wu-zhong Guo, Song He, Yan Liu, and Jie-qiang Wu for insightful discussions and constructive comments.
@@ Appendix A Analytical continuation
$\hskip-28.45274ptG_{8}(p)=\sum_{a_{1}<a_{2},b_{1}<b_{2}}\bigg[\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi b_{12}}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{1})}{2p}\sin\frac{\pi(a_{2}-b_{2})}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{2})}{2p}\sin\frac{\pi(a_{2}-b_{1})}{2p}\big)^{2}}\bigg],\$

### v2

@@ 3.3 Case III
No higher-order analytical results exist in the literature for these cases, so we compare our predictions with numerical results in the spin-1/2 XX chain. In the XX chain, the above CFT results are expected to apply in the limit $1\ll\ell\ll L$. To match CFT and spin chain results, we require high-efficiency and high-precision numerical evaluation of subsystem fidelity in the spin chain. We calculate subsystem fidelity in the XX chain using the truncation method with truncation number $t=10$ from [48] and show results in Figure 1. We define the subtracted fidelity
@@
which is removed from both CFT and spin chain results for clearer comparison. The figure uses $(\alpha,\bar{\alpha})=(1,0)$, showing good agreement between CFT predictions and spin chain results. For other values of $(\alpha,\bar{\alpha})$, the agreement with theory is not always consistent, likely due to finite-size effects or numerical limitations. Additional results are shown in Figure 2. Although the last column does not always show perfect agreement, a clear trend emerges where, with the increase of the truncation number $t$, the numerical results approach the analytical short-interval prediction for $1\ll\ell\ll L$.
Figure 1: Comparison between CFT predictions (solid red lines) and $X$ spin chain results (symbols) for the subtracted fidelity $F_{(2)}-F$. In all panels, the spin chain results are obtained from the truncation method with truncation number $t=10$. The good agreement validates the short-interval expansion approach.
Figure 2: We show additional comparisons of the subtracted fidelity $|F_{(2)}-F|$ between CFT predictions (solid red lines) and $X$ spin chain results (symbols). The four columns correspond to truncation numbers $t=4,6,8,10$, demonstrating that the numerical results converge to the analytical predictions as $t$ increases.
3.4 Validity range
The short-interval expansion derived from the OPE of twist operators is expected to be valid only for $\frac{\ell}{L}\ll 1$. A key question is to determine the range of convergence for this expansion and the validity of its finite truncation. This is challenging due to the currently limited number of analytical terms and the precision of our numerical data. Nevertheless, we can analyze this for the known exact analytical results. To this end, we employ the exact fidelity (3.4) and its subtracted fidelity defined as
$F_{(2)}\equiv 1-\frac{[(\alpha-\alpha^{\prime})^{2}+(\bar{\alpha}-\bar{\alpha}^{\prime})^{2}]\pi^{2}x^{2}}{16}.$ (3.9)
Figure 3 shows the error of the relative subtracted fidelity, $|1-F_{(2)}/F|$, revealing distinct ranges of validity for different fidelities.
Figure 3: A comparison of the exact fidelity $F(\rho_{A,\alpha,\bar{\alpha}},\rho_{A,\alpha^{\prime},\bar{\alpha}^{\prime}})$ (solid lines, (3.4)) and the subtracted fidelity (red dashed lines, (3.9)) is shown on the left, with the corresponding relative error on the right. The line colors in both panels represent different values of $(\alpha-\alpha^{\prime})^{2}+(\bar{\alpha}-\bar{\alpha}^{\prime})^{2}$, as shown in the legend.
@@ 4.1 Case I
Using the same formula (3.1) as in the bosonic case, we reproduce the short interval expansion of these results.
@@ 4.2 Case II
From the contributions $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}$ (2.30) with $\mathcal{X}=\varepsilon,T,\bar{T}$, $\mathcal{F}^{(1/2)}_{T\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=\varepsilon$, $\mathcal{F}^{(1/2)}_{\bar{T}\mathcal{X}\mathcal{X}}$ (2.5) with $\mathcal{X}=\varepsilon$, and $\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon\varepsilon\varepsilon}$ (2.48), and using the expansion
$F(\rho_{A},\sigma_{A})=1+\ell^{2}\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon}+\ell^{4}\big(\mathcal{F}^{(1/2)}_{\varepsilon\varepsilon\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{T\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{\bar{T}\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}_{TT}+\mathcal{F}^{(1/2)}_{\bar{T}\bar{T}}\big)+O(\ell^{6}),$ (4.4)
@@ 4.3 Case III
We compare these predictions with numerical results in the critical Ising chain, obtained using the truncation method from [48] with truncation number $t=10$, as shown in Figure 4. We define the subtracted fidelity
@@
In this section, we have only included the contributions from the identity conformal family. If the lightest non-identity operator has conformal weight $\Delta$, according to (2.21) its leading contribution to the short interval expansion of the fidelity would be of order $O(\ell^{2\Delta})$. For the results (5.2) and (5.3) to be valid, we need $\Delta>3$, and for the result (5.5) to be valid, we need $\Delta>4$.
The fidelity (5.5) quantifies how corrections from a quantum theory of gravity encode distinctions between microscopic states into the data accessible to small subsystems [52, 53]. The result directly resolves a puzzle concerning the perturbative distinguishability of black hole microstates. This puzzle, noted in footnote 17 of [53], stemmed from a tension: studies of fixed-area states, superpositions of energy eigenstates, suggested that distinguishing microstates required non-perturbative precision, whereas evidence from the $1/c$ expansion in holographic CFTs indicated a perturbative effect [54, 30, 32, 33, 34, 55]. The ambiguity arose because the distinguishability of fixed-area states depends on the details of the superposition. Our analysis resolves this by proving that for exact primary states, genuine energy eigenstates dual to black hole microstates, the subsystem trace distance exhibits a universal $1/c$ scaling. This firmly establishes that quantum gravity corrections encode perturbative distinctions between a microstate and the thermal state into small, accessible subsystems.
This demonstrated distinguishability challenges the standard eigenstate thermalization hypothesis (ETH) [56, 57], which posits that individual high-energy eigenstates are locally indistinguishable from a thermal ensemble. The violation of ETH in these holographic systems necessitates a generalized description of thermalization. A compelling framework is a generalized Gibbs ensemble (GGE) [58, 59] incorporating Korteweg-de Vries (KdV) charges [60, 59, 61], which accounts for additional conserved quantities that distinguish microstates of the same energy. This picture is supported by independent evidence from identity conformal family operators [62] and large subsystem Rényi entropy [63, 64].
Thus, our proof of perturbative distinguishability for exact eigenstates resolves the foundational puzzle and challenges the standard ETH. While our results necessitate a generalized thermalization picture, the proposal that a GGE with KdV charges universally describes all local observables remains a highly motivated conjecture for future work.
@@ Acknowledgements
J.Z. is grateful to the organizers and participants of the Annual Academic Conference of the Gravitation and Relativistic Astrophysics Division of the Chinese Physical Society (Kunming, April 2025), the Symposium on Gravity and Cosmology (Xi’an, June 2025), the Symposium on Strings, Gravity, and Gravitational Waves (Rizhao, August 2025), and the Symposium on Gravity and Cosmology (Hefei, November 2025), where preliminary results of this work were presented and benefitted from valuable discussions. J.Z. also wishes to thank the Institute of Fundamental Physics and Quantum Technology at Ningbo University for its hospitality during a productive visit where part of this research was presented, discussed and completed. In addition, J.Z. thanks Li-Ming Cao, Bin Chen, Anatoly Dymarsky, Wu-zhong Guo, Song He, Yan Liu, Jian-Xin Lu, Jia-Rui Sun, Shao-Jiang Wang, Jie-qiang Wu, and Li-Xin Xu for insightful discussions and constructive comments.
@@ Appendix A Analytical continuation
There is an alternative way to derive the analytical continuation (A.5). Using results in [26], we write $G_{1}(p,\Delta)$ in (A.1) as
$G_{1}(p,\Delta)=\frac{p(p-1)}{2}g(0)+p\sum_{k=1}^{+\infty}[pg(pk)-g(k)],$ (A.6)
with the definition
$g(k)=\frac{2^{2\Delta}}{\pi\cos(\pi\Delta)}\sin[\pi(\Delta-k)]\int_{0}^{\pi/2}(\sin\varphi)^{2(\Delta+k)-1}(\cos\varphi)^{2(\Delta-k)-1}\mathrm{d}\varphi.$ (A.7)
Thus we have
$G_{1}\Big(\frac{1}{2},\Delta\Big)=-\frac{1}{8}g(0)+\frac{1}{2}\sum_{k=1}^{+\infty}\Big[\frac{1}{2}g\Big(\frac{k}{2}\Big)-g(k)\Big],$ (A.8)
Using
$\sum_{k=1}^{+\infty}\Big\{\frac{1}{2}\sin\Big[\pi\Big(\Delta-\frac{k}{2}\Big)\Big](\tan\varphi)^{k}-\sin[\pi(\Delta-k)](\tan\varphi)^{2k}\Big\}=-\frac{1}{2}\sin\varphi\cos(\pi\Delta+\varphi),$ (A.9)
we get
$G_{1}\Big(\frac{1}{2},\Delta\Big)=\frac{2^{2\Delta}}{\pi\cos(\pi\Delta)}\int_{0}^{\pi/2}\Big[-\frac{1}{8}\sin(\pi\Delta)(\sin\varphi)^{2\Delta-1}(\cos\varphi)^{2\Delta-1}$ (A.10)
[…]
From
$\int_{0}^{\pi/2}\sin^{\alpha}\varphi\cos^{\beta}\varphi\mathrm{d}\varphi=\frac{\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{\beta+1}{2})}{2\Gamma(\frac{\alpha+\beta}{2}+1)},$ (A.11)
which is convergent for $\alpha>-1$, $\beta>-1$, we finally obtain the same analytical continuation (A.5). This verifies the method in the previous paragraph. The method in this paragraph relies on special transformations of the expressions, while the method in the previous paragraph is simpler and easier to generalize to more complicated cases, and in the following we will use the simpler method to do the analytical continuations.
@@
$\hskip-28.45274ptG_{8}(p)=\sum_{a_{1}<a_{2},b_{1}<b_{2}}\bigg[\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi b_{12}}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{1})}{2p}\sin\frac{\pi(a_{2}-b_{2})}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{2})}{2p}\sin\frac{\pi(a_{2}-b_{1})}{2p}\big)^{2}}\bigg],$